OpenStudy (anonymous):

2/x + 3/x-3 =1 Solve?? *sad face*

OpenStudy (anonymous):

3/x-3 means 3/(x-3)?

OpenStudy (anonymous):

or -3+3/x ?

OpenStudy (anonymous):

the x-3 is the denominator

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2/x + 3/(x-3) =1 2x(x-3)/x+3x(x-3)/(x-3) = x(x-3) 2(x-3)+3x=x(x-3) 2x-6+3x=x^2-3x x^2-8x+6=0 x^2-8x+6+10=10 x^2-8x+16=10 (x-4)^2 = 10 x-4 = √10 or x-4 = -√10 x= √10 +4 or x= -√10 +4

OpenStudy (anonymous):

ok, thanks, wait a sec! i have three more, Please help me, i have only minutes left

OpenStudy (anonymous):

\[x-2\sqrt{2x}=16\]

OpenStudy (anonymous):

x-2√2√x-16=0 (√x)^2-2√2√x-16=0 (√x)^2-2√2√x+2-16=2 (√x-√2)^2=18 √x-√2=3√2 or √x-√2=-3√2 √x=4√2 or √x=-2√2 √x should be always positive. so √x = 4√2 x = 32

OpenStudy (anonymous):

:D thanks, ill be right back

OpenStudy (anonymous):

x-2sqrt5x=15

OpenStudy (anonymous):

\[x-2\sqrt{5x}=15\]

OpenStudy (anonymous):

(√x)^2-2√5√x+5=15+5 (√x-√5)^2=20 √x-√5=2√5 or √x-√5=-2√5 √x=3√5 or √x=-√5 √x should be positive. so, x = 45

OpenStudy (anonymous):

okay thanks, can you help me with one more? i have three more left, but one should be enough :)